# SEA-SQL: Semantic-Enhanced Text-to-SQL with Adaptive Refinement

### v1

@@ 4.3 Dynamic Execution Adjustment
(a) The process of Dynamic Execution Adjustment.
(b) An example of Dynamic Execution Adjustment.
Figure 3: Dynamic Execution Adjustment: Process and Example.
The primary focus of the Task-to-SQL task is to guarantee the executability of the generated SQL. However, a significant portion of prior research concentrates on the refinement of SQL exclusively through the LLM, ignoring whether the execution is successful or not (Pourreza and Rafiei, 2023). As depicted in Figure 3(a), we propose dynamic execution adjustment, which corrects unexecutable SQL queries by leveraging the LLM to both identify the error reason in the SQL query and regenerate it for execution. Inspired by Chain-of-Thought (CoT) (Wei et al., 2022) and REACT (Yao et al., 2023), we divide the adjustment process into three steps: execute, reflect, and correct. Through external execution of SQL API calls and continuous internal reflection and correction within the LLM, we ensure that the SQL query eventually can be successfully executed.
@@
Detailed zero-shot prompts of these steps can be found in Appendix C. Figure 3(b) illustrates an example of dynamic execution adjustment process, the LLM dynamically corrects errors in SQL queries by utilizing error feedback and its reflection information. When an error is detected by execution, the LLM conducts a thorough reflection of its error reason and try to correct it, leading to the generation of a refined SQL query. This refined query is then undergoes another execution attempt. If additional errors emerge during this iterative process, the LLM methodically reflects upon and corrects the SQL iteratively until a version capable of successful execution is obtained.
@@ 5.1 Experiment Setup
Datasets. The Spider dataset, as described in (Yu et al., 2018b), is a comprehensive semantic parsing and Text-to-SQL dataset of significant scale and complexity. This dataset comprises 10,181 questions and 5,693 unique complex SQL queries. It covers 200 datasets featuring multiple tables and spans 138 diverse domains.
@@
The BIRD dataset, introduced in (Li et al., 2023b), serves as a substantial benchmark for large-scale database tasks related to Text-to-SQL. This dataset includes 12,751 pairs of Text-to-SQL data and encompasses 95 databases, with a collective size of 33.4 GB. Covering 37 professional domains, BIRD stands out as the first Text-to-SQL benchmark to emphasize efficiency, encouraging the development of more efficient query methods tailored for massive and noisy database contents.
@@
To ensure the robustness of our model, we conduct a thorough evaluation using another distinct dataset: Spider-Realistic (Deng et al., 2021) consisting of 508 samples. This dataset is curated from the Spider dataset, wherein questions are strategically modified to emulate real-world application scenarios. Specifically, Spider-Realistic omits explicitly mentioned column names in questions.
@@
Evaluation Metrics. To evaluate the performance of the Text-to-SQL task, following (Zhong et al., 2020) and (Li et al., 2023b), we adopt two metrics: Execution accuracy (EX) and Valid Efficiency Score (VES). We use EX for Spider, and both EX and VES for BIRD.
  • The execution accuracy goes beyond syntax and compares whether the result obtained from executing the predicted SQL query is identical to the result of the gold SQL query. In the BIRD dataset, ACC is equal to EX.
  • The valid efficiency score is designed to holistically evaluate a model’s performance by considering both the efficiency and accuracy of the execution results of valid SQL queries. It provides a comprehensive measure that takes into account the quality and speed of the model’s output.
@@
LLMs. We employ two distinct LLMs for our task: GPT-3.5, a black-box model accessible solely through the API, and Mistral-7B, an open-source model that allows us not only to utilize it via the API but also affords the flexibility to freely fine-tune its parameters.
@@
Implementation Details. We employ GPT-3.5  [^1] We use gpt-3.5-turbo-16k-0613 to conduct our experiment. for SQL query generation and dynamic execution adjustment. To mitigate biases in GPT-3.5, we utilize a fine-tuned Mistral-7B model as our adaptive bias eliminator. The Mistral-7B model is fine-tuned using LoRA (Xu et al., 2023), deepspeed (Aminabadi et al., 2022) and flash attention (Dao et al., 2022), employing a batch size of 64 and a learning rate of 2e-4. The training is conducted on the Spider and BIRD training datasets for one epoch using 8 × 32G V100 GPUs in 1.5 hours. Additionally, we incorporate a linear warm-up with 10% of the training steps and employ linear decay to adjust the learning rate during training.
@@ 5.2 Overall Performance
method model ZS FS FT dev
DIN-SQL (Pourreza and Rafiei, 2023) GPT-4 82.8
DAIL-SQL (Gao et al., 2024) GPT-4 83.5
MAC-SQL (Wang et al., 2023a) GPT-4 86.8
[…]
DIN-SQL (Pourreza and Rafiei, 2023) GPT-3.5 74.9
DAIL-SQL (Gao et al., 2024) GPT-3.5 78.2
MAC-SQL (Wang et al., 2023a) GPT-3.5 80.6
[…]
T5-3B + PICARD (Scholak et al., 2021) T5-3B 79.9
Supervised Fine-Tuning (Gao et al., 2024) Llama2-7b 66.7
Supervised Fine-Tuning (Gao et al., 2024) Llama2-13b 67.0
SEA-SQL(ours) GPT-3.5 + Mistral-7B 83.6
Table 1: SEA-SQL and baselines on Spider’s development set (EX result).
Spider Result. Table 1 presents the performance comparison between our proposed framework and baseline approaches on the Spider dataset’s development set. Notably, it’s essential to mention that Spider has closed its test set for evaluation. Leveraging GPT-3.5, SEA-SQL outperforms other methods that either use GPT-3.5 or are fine-tuned (FT), achieving the a high EX score on the development set.
A distinguishing factor of our approach is the use of zero-shot (ZS) prompts, as opposed to the few-shot (FS) prompts that is predominantly used in other methods. This highlights the simplicity and effectiveness of our approach.
Compared to fine-tuned approaches (Gao et al., 2024), the utilization of a small LLM as a bias eliminator is more effective than directly employing a small LLM to generate SQL. In comparison to GPT-3.5 based approaches, SEA-SQL significantly outperforms them, emphasizing the effectiveness of combining GPT-3.5 with fine-tuning as a promising direction for further exploration. Moreover, the results remain competitive with DIN-SQL and DAIL-SQL in the scenario of GPT-4, demonstrating SOTA performance based on GPT-3.5 and maintaining competitiveness relative to GPT-4.
method model dev test
ACC VES ACC VES
@@
SEA-SQL GPT-3.5 + 56.13 60.13 53.66 57.00
(ours) Mistral-7B
Table 2: SEA-SQL and baselines on BIRD (ACC is equal to EX result).
BIRD Result. Table 2 showcases the performance of our framework on the BIRD dataset. SEA-SQL demonstrates substantial performance superiority over other GPT-3.5-based methods. Particularly noteworthy is its competitiveness with DIN-SQL and DAIL-SQL on the development set. Our approach surpasses the current best GPT-3.5-based MAC-SQL ACC result by a significant margin of 11.02% on the development set, thereby establishing a new SOTA performance based on GPT-3.5.
@@ 5.3 Ablation Study
method Simple Mod. Chall. All
SEA-SQL 64.11 47.74 31.94 56.13
@@
w/o DEA 62.27 43.44 27.28 53.32
w/o ABE & DEA 59.57 37.85 21.53 49.41
w/o SES 37.84 18.06 15.97 29.79
Table 3: The EX results of ablation studies on the BIRD development set.
We conducted a series of ablations to investigate the pivotal components of our framework. Four sets of ablation experiments were performed on the BIRD development set, each targeting a specific aspect of our approach: the adaptive bias elimination (ABE), dynamic execution adjustment (DEA), both the ABE and DEA, and the semantic-enhanced schema (SES), which was replaced with standard schema.
Table 3 presents the experimental results of the ablation study. Additionally, it illustrates the model’s performance across varying complexities of SQL problems, categorized as Simple, Moderate (Mod.), and Challenging (Chall.).
The results of the ablation study demonstrate the significance of both the bias eliminator and dynamic execution adjustment components. Removal of either of these elements resulted in a considerable decline in performance. With the simultaneous removal of both leading to a substantial deterioration in results. Notably, once the semantic-enhanced schema is replaced with the standard schema, the final results decrease significantly. This indicates that including the values of columns related to the question facilitates a more comprehensive understanding of the database information for the LLM, thus leads to the generation of superior SQL queries.
Overall, the results definitely underscore the crucial role played by each component. Notably, the absence of any module consistently resulted in a noticeable decrease in overall performance. This highlights the interdependence of the modules within our framework and underscores their collective significance in attaining optimal results. Appendix A displays results of the experiment assessing the impact of different sample values.
@@ 5.4 Cost Analysis
One primary concern in the development of Text-to-SQL technology is its accessibility to ordinary users. To address this concern, we conducted an evaluation of our approach, comparing it against established baselines in terms of cost effectiveness (up to May 2024)  [^2] We use gpt-3.5-turbo-16k-0613 to calculate the cost of GPT-3.5 and gpt-4(8k) to calculate the cost of GPT-4., as outlined in Table 4, where "pre" denotes the cost of pre-processing, "generate" refers to the cost required to generate each SQL query, and "all" represents the total cost when generating all SQL queries. Our findings reveal a significant cost advantage with our framework over the GPT-4 based approach. Although our GPT-3.5 model incurs slightly higher costs with other GPT-3.5 based methods, primarily attributed to pre-processing, the expense is significantly mitigated after training the bias eliminator. Notably, the cost of generating a single SQL query is still minimal, ranging from 9% to 58% compared to other GPT-3.5 based methods due to zero-shot prompts.
method model pre generate all
DIN-SQL GPT-4 - 0.7867 1207
DAIL-SQL GPT-4 - 0.1232 189
MAC-SQL GPT-4 - 0.2151 330
[…]
DAIL-SQL GPT-3.5 - 0.0113 17
MAC-SQL GPT-3.5 - 0.0209 32
SEA-SQL GPT-3.5 44.78 0.0065 55
Table 4: The cost (\$) of different methods on the BIRD development set varies.
5.5 Robustness

### v2

@@ 4.3 Dynamic Execution Adjustment
Figure 4: The process of Dynamic Execution Adjustment.
Figure 5: An example of Dynamic Execution Adjustment.
Although SQL queries may be free of biases, they can still fail to execute due to issues such as column mismatches or runtime errors. These issues are not directly related to bias, but they determine whether the SQL can ultimately execute successfully, which is a crucial aspect of the Text-to-SQL task. However, a significant portion of prior research concentrates on the refinement of SQL exclusively through the LLM, ignoring whether the execution is successful or not Pourreza and Rafiei (2023). As depicted in Figure 4, we propose dynamic execution adjustment, which corrects unexecutable SQL queries by leveraging the LLM to both identify the error reason in the SQL query and regenerate it for execution. Inspired by Chain-of-Thought (CoT) Wei et al. (2024) and REACT Yao et al. (2023), we divide the adjustment process into three steps: execute, reflect, and correct. Through external execution of SQL API calls and continuous internal reflection and correction within the LLM, we ensure that the SQL query eventually can be successfully executed.
@@
Detailed zero-shot prompts of these steps can be found in Appendix C. Figure 5 illustrates an example of dynamic execution adjustment process, the LLM dynamically corrects errors in SQL queries by utilizing error feedback and its reflection information. When an error is detected by execution, the LLM conducts a thorough reflection of its error reason and try to correct it, leading to the generation of a refined SQL query. This refined query is then undergoes another execution attempt. If additional errors emerge during this iterative process, the LLM methodically reflects upon and corrects the SQL iteratively until a version capable of successful execution is obtained.
@@ 5.1 Experiment Setup
Datasets. The Spider dataset, as described in Yu et al. (2018b), is a comprehensive semantic parsing and Text-to-SQL dataset of significant scale and complexity. This dataset comprises 10,181 questions and 5,693 unique complex SQL queries. It covers 200 datasets featuring multiple tables and spans 138 diverse domains. It comprises 8,659 instances for the training set and 1,034 instances for the development set.
@@
The BIRD dataset, introduced in Li et al. (2023c), serves as a substantial benchmark for large-scale database tasks related to Text-to-SQL. This dataset includes 12,751 pairs of Text-to-SQL data and encompasses 95 databases, with a collective size of 33.4 GB. Covering 37 professional domains, BIRD stands out as the first Text-to-SQL benchmark to emphasize efficiency, encouraging the development of more efficient query methods tailored for massive and noisy database contents. It comprises 9,428 instances for the training set, 1,534 instances for the development set, and 1,789 instances for the concealed test set.
@@
To ensure the robustness of our model, we conduct a thorough evaluation using another distinct dataset: Spider-Realistic Deng et al. (2021) consisting of 508 samples. This dataset is curated from the Spider dataset, wherein questions are strategically modified to emulate real-world application scenarios. Specifically, Spider-Realistic omits explicitly mentioned column names in questions. It comprises 508 instances for the development set.
@@
Evaluation Metrics. To evaluate the performance of the Text-to-SQL task, following Zhong et al. (2020) and Li et al. (2023c), we adopt two metrics: Execution Accuracy (EX) and Valid Efficiency Score (VES).
  • The Execution Accuracy (EX) is the proportion of examples in the evaluation dataset where the execution results of the predicted SQL queries are the same as those of the gold SQL queries. For an execution result $Y_{n}$ of a gold SQL denoted by $V_{n}$, and an execution result of a predicted SQL $\hat{Y}_{n}$ denoted by $\hat{V}_{n}$, EX is computed using the formula:
    $\text{EX}=\frac{\sum_{n=1}^{N}\Phi(V_{n},\hat{V}_{n})}{N}$
    where $N$ represents the total number of SQL queries, and the indicator function $\Phi(\cdot)$ is defined as:
    $\Phi(V,\hat{V})=\begin{cases}1,&\text{if }V=\hat{V}\\ 0,&\text{if }V\neq\hat{V}\end{cases}$
  • The Valid Efficiency Score (VES) could assess the efficiency of valid SQL queries generated by models. A "valid SQL" is defined as a predicted SQL query whose execution result matches its corresponding gold SQL result. Queries that fail to generate correct execution results are deemed invalid as they are ineffective for user needs, irrespective of efficiency. Therefore, the VES metric takes into account both the effectiveness and the efficiency of execution, providing a comprehensive measure of model performance. The VES can be calculated as follows:
    $\text{VES}=\frac{\sum_{n=1}^{N}\Phi(V_{n},\hat{V}_{n})\cdot R(Y_{n},\hat{Y}_{n})}{N}$
    where $R(\cdot)$ evaluates the relative execution efficiency of the predicted SQL in relation to the gold SQL:
    $R(Y_{n},\hat{Y}_{n})=\sqrt{\frac{E(Y_{n})}{E(\hat{Y}_{n})}}$
    here $E(\cdot)$ is a function that assesses the execution efficiency of each SQL in a specific setting Li et al. (2023c).
Method Model Zero-Shot Few-Shot Fine-Tune dev
GPT-4 Pourreza and Rafiei (2023) GPT-4 72.9
DIN-SQL Pourreza and Rafiei (2023) GPT-4 82.8
DAIL-SQL Gao et al. (2024) GPT-4 83.5
MAC-SQL Wang et al. (2023a) GPT-4 86.8
[…]
DIN-SQL Pourreza and Rafiei (2023) GPT-3.5 74.9
DAIL-SQL Gao et al. (2024) GPT-3.5 78.2
MAC-SQL Wang et al. (2023a) GPT-3.5 80.6
[…]
T5-3B + PICARD Scholak et al. (2021) T5-3B 79.9
Supervised Fine-Tuning Gao et al. (2024) Llama2-7b 66.7
Supervised Fine-Tuning Gao et al. (2024) Llama2-13b 67.0
SEA-SQL(ours) GPT-3.5 + Mistral-7B 83.6
Table 1: SEA-SQL and baselines on Spider development set (EX result), where the best outcomes are highlighted in bold, while the second highest results are underlined.
@@
LLMs. We employ two distinct LLMs for our task: GPT-3.5, a black-box model accessible solely through the API, and Mistral-7B, an open-source model that allows us not only to utilize it via the API but also affords the flexibility to freely fine-tune its parameters. We employ GPT-3.5  [^1] We use gpt-3.5-turbo-16k-0613 to conduct our experiment. for SQL query generation and dynamic execution adjustment. To mitigate biases in GPT-3.5, we utilize a fine-tuned Mistral-7B model as our adaptive bias eliminator.
@@
Implementation Details. The Mistral-7B model is fine-tuned using LoRA Xu et al. (2023), deepspeed Aminabadi et al. (2022) and flash attention Dao et al. (2022), employing a batch size of 64 and a learning rate of 2e-4. The training is conducted on the Spider and BIRD training datasets for one epoch using 8 × 32G V100 GPUs in 1.5 hours. Additionally, we incorporate a linear warm-up with 10% of the training steps and employ linear decay to adjust the learning rate during training. For semantic-enhanced schema, we retrieve the top 1000 sample values and retain the top 2 after "bridging". To prevent the system from entering an infinite loop, we have capped the maximum iterations for dynamic execution adjustment at 10.
Baselines. We conducted experiments on both the BIRD and Spider datasets, comparing our method with the following baselines:
  • GPT-4 Pourreza and Rafiei (2023) utilizes a straightforward zero-shot Text-to-SQL prompt for SQL generation.
  • C3 Dong et al. (2023) employs clear prompting, calibration with hints, and ensures consistent output through self-consistency to generate SQL queries.
  • DIN-SQL Pourreza and Rafiei (2023) decomposes the generation task into smaller sub-problems. The solutions to these sub-problems are then inputted into LLMs to produce the final SQL output.
  • DAIL-SQL Gao et al. (2024) leverages few-shot prompting by retrieving examples from the training set to facilitate the generation of SQL queries.
  • MAC-SQL Wang et al. (2023a) completes the text-to-SQL task through a multi-agent collaboration approach, involving schema selection, question decomposition, and SQL refinement.
  • T5-3B Scholak et al. (2021) fine-tunes the T5-3B model for text-to-SQL tasks.
  • T5-3B + PICARD Scholak et al. (2021) fine-tunes the T5-3B model for text-to-SQL tasks, with PICARD applying syntax rules to ensure the generated SQL is parsed incrementally, verifying each step against predefined syntax and structural constraints.
  • Supervised Fine-Tuning Gao et al. (2024) involves directly fine-tuning LLMs specifically for the text-to-SQL task.
@@ 5.2 Overall Performance
Method Model Zero-Shot Few-Shot Fine-Tune dev test
EX VES EX VES
GPT-4 GPT-4 46.35 49.77 54.89 60.77
@@
SEA-SQL(ours) GPT-3.5 + Mistral-7B 56.13 60.83 53.66 57.00
Table 2: SEA-SQL and baselines on BIRD.
Method BIRD Spider
EX VES
Simple Mod. Chall. All Simple Mod. Chall. All EX
SEA-SQL 64.11 47.74 31.94 56.13 69.63 51.71 33.72 60.83 83.6
[…]
   w/o DEA 62.27 43.44 27.28 53.32 67.80 46.72 29.64 57.83 83.6
   w/o ABE & DEA 59.57 37.85 21.53 49.41 63.98 41.03 22.50 53.13 80.9
Table 3: The results of ablation studies on the development set.
Values Num BIRD Spider
EX VES
Simple Mod. Chall. All Simple Mod. Chall. All EX
zero 49.51 30.11 21.53 41.00 55.65 30.20 22.13 44.79 76.4
one 59.24 36.34 22.22 48.83 63.55 38.47 22.53 52.10 79.5
two 59.57 37.85 21.53 49.41 63.98 41.03 22.50 53.13 80.9
three 59.14 36.99 20.83 48.83 61.97 40.88 23.29 51.95 78.5
Table 4: The results of the number of different sample values on the development set.
Spider Result. Table 1 presents the performance comparison between our proposed framework and baseline approaches on the Spider dataset’s development set. Notably, it’s essential to mention that Spider has closed its test set for evaluation. Leveraging GPT-3.5, SEA-SQL outperforms other methods that either use GPT-3.5 or are fine-tuned, achieving a high EX score on the development set.
A distinguishing factor of our approach is the use of zero-shot prompts, as opposed to the few-shot prompts that is predominantly used in other methods. This highlights the simplicity and effectiveness of our approach.
Compared to fine-tuned approaches Gao et al. (2024), the utilization of a small LLM as a bias eliminator is more effective than directly employing a small LLM to generate SQL. In comparison to GPT-3.5 based approaches, SEA-SQL significantly outperforms them, emphasizing the effectiveness of combining GPT-3.5 with fine-tuning as a promising direction for further exploration. Moreover, the results remain competitive with DIN-SQL and DAIL-SQL in the scenario of GPT-4, demonstrating SOTA performance based on GPT-3.5 and maintaining competitiveness relative to GPT-4.
BIRD Result. Table 2 showcases the performance of our framework on the BIRD dataset. SEA-SQL demonstrates substantial performance superiority over other GPT-3.5-based methods. Particularly noteworthy is its competitiveness with DIN-SQL and DAIL-SQL on the development set. Our approach surpasses the current best GPT-3.5-based MAC-SQL EX result by a significant margin of 11.02% on the development set, thereby establishing a new SOTA performance based on GPT-3.5.
@@ 5.3 Ablation Study
We conducted a series of ablations to investigate the pivotal components of our framework. Four sets of ablation experiments were performed on the BIRD development set, each targeting a specific aspect of our approach: the adaptive bias elimination (ABE), dynamic execution adjustment (DEA), and both the ABE and DEA.
Table 3 presents the experimental results of the ablation study. Additionally, it illustrates the model’s performance across varying complexities of SQL problems, categorized as Simple, Moderate (Mod.), and Challenging (Chall.). These categories are defined by human experts and consider factors such as Question Understanding, Knowledge Reasoning, Data Complexity, and SQL Complexity Li et al. (2023c).
The ablation study on the BIRD dataset underscores the importance of both the adaptive bias elimination and the dynamic execution adjustment components in handling SQL queries of varying complexity. Removing either component significantly reduced performance, and eliminating both drastically worsened the outcomes. The impact of removing these components was particularly severe for moderate and complex SQL queries, while simpler SQLs experienced a smaller decline in performance. This suggests that our proposed SEA-SQL framework is particularly beneficial for generating more complex SQL queries. In the case of the Spider dataset, the DEA component did not influence the final results due to the dataset’s relative simplicity. All generated bias-eliminated SQL queries were executable, rendering the dynamic execution adjustment superfluous in improving performance.
Method Model Pre (\$) Generate / Query (\$) All (\$) GPU-Time / Query (s) API-Time / Query (s) Full-Time / Query (s)
GPT-4 GPT-4 - 0.0351 54 - 1.17 1.17
DIN-SQL GPT-4 - 0.7867 1207 - 13.86 13.86
DAIL-SQL GPT-4 - 0.1232 189 - 12.29 12.29
MAC-SQL GPT-4 - 0.2151 330 - 15.84 15.84
@@
DAIL-SQL GPT-3.5 - 0.0113 17 - 12.29 12.29
MAC-SQL GPT-3.5 - 0.0209 32 - 15.84 15.84
SEA-SQL GPT-3.5 44.78 0.0065 55 1.09 7.37 8.46
Table 5: The cost of different methods on the BIRD development set.
Overall, the results definitely underscore the crucial role played by each component. Notably, the absence of any module consistently resulted in a noticeable decrease in overall performance. This highlights the interdependence of the modules within our framework and underscores their collective significance in attaining optimal results.
5.4 Results on Different Sample values
To investigate the impact of the semantic-enhanced schema on performance, we conducted comparative analyses between the standard schema and the semantic-enhanced schema, utilizing varying numbers of sample values. Given that modifying the number of samples necessitates retraining of the adaptive bias eliminator, our evaluation focused solely on the performance of pre-SQL. Due to the excessive number of sample values, the input may become too lengthy, making it challenging for the LLM to accurately comprehend the database information. Consequently, we conducted tests using 0, 1, 2, and 3 samples, as illustrated in the Table 4. The results provided significant insights into the importance of sample values in the effectiveness of our methodology.
When the number of values is zero, essentially using standard schema, a notable 17% decrease is observed. This reduction underscores the significance of our semantic-enhanced schema in achieving optimal results.
While using only two sample values yielded the best results exclusively for the moderate complexity level of the BIRD dataset, this approach closely approached optimal performance for both the simple and complex complexity levels. Additionally, the average performance achieved by using two sample values surpassed that obtained with either one or three sample values across both the BIRD and Spider datasets. This improvement likely occurs because two values offer more informative content than a single value, while three values can make the input overly long, potentially degrading performance.
@@ 5.5 Cost Analysis
One primary concern in the development of Text-to-SQL technology is its accessibility to ordinary users Jiang et al. (2024); Jin et al. (2024); Li et al. (2023a). To address this concern, we conducted an evaluation of our approach, comparing it against established baselines in terms of cost effectiveness (up to May 2024)  [^2] We use gpt-3.5-turbo-16k-0613 to calculate the cost of GPT-3.5 and gpt-4(8k) to calculate the cost of GPT-4., as outlined in Table 5, where "Pre" denotes the cost of pre-processing, "Generate" refers to the cost required to generate each SQL query, and "All" represents the total cost when generating all SQL queries. "GPU-Time" denotes the duration each query utilizes GPU, "API-Time" refers to the API request time for each SQL query, and "Full-Time" indicates the total time required to generate a complete SQL query. Our findings reveal a significant cost advantage with our framework over the GPT-4 based approach. Since SEQ-SQL only uses 25GB of GPU and takes 1.09 seconds to generate a single SQL, this generation overhead can be considered negligible. Although our GPT-3.5 model incurs slightly higher costs with other GPT-3.5 based methods, primarily attributed to pre-processing, the expense is significantly mitigated after training the bias eliminator. Notably, the cost of generating a single SQL query is still minimal, ranging from 9% to 58% compared to other GPT-3.5 based methods due to zero-shot prompts. Although our method involves executing SQL queries, it is more efficient than other approaches. The use of a zero-shot prompt significantly reduces the overall execution time by minimizing the prompt length compared to few-shot methods.
@@
5.6 Robustness
Generation Model Bias Elimination Model EX VES
Simple Mod. Chall. All Simple Mod. Chall. All
GPT-3.5 Mistral-7B 64.11 47.74 31.94 56.13 69.63 51.71 33.72 60.83
GPT-4o-mini Mistral-7B 65.30 44.73 36.11 56.32 68.52 46.88 36.70 58.98
Qwen-2.5-72b-it Mistral-7B 64.86 49.46 35.42 57.43 68.22 58.20 38.04 60.53
Llama-3.1-70b-it Mistral-7B 66.70 53.76 47.22 60.95 70.45 62.30 51.41 66.19
GPT-3.5 Gemma2-7B 62.70 48.17 33.33 55.54 68.05 51.78 33.54 59.88
GPT-3.5 Llama2-7B 62.59 43.66 31.25 53.91 67.19 47.10 32.20 57.82
Table 7: The results of different backbones on the BIRD development set.